OpenStudy (anonymous):

write an equation in point slope format of the line that passes through the given point and with the given slope 'm' (3,-4);m=6 (-2,-7);m=4/5

OpenStudy (nenadmatematika):

y+4=6(x-3) and y+7=(4/5)(x+2)

OpenStudy (anonymous):

Thank you :D